OpenStudy (anonymous):

what is another way of writing a=logbase4 ^16

OpenStudy (anonymous):

16^a=4

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

\[a=\log_{4}16 \]\[a^4=16\]\[a=2^4\]

OpenStudy (anonymous):

No. Base stays base.

OpenStudy (anonymous):

\[4^{a}=16\]

OpenStudy (anonymous):

I see log base 4 of 16 as asking the question... How many powers of 4 are in 16. It would be 4 squared is sixteen, so a is 2.

OpenStudy (anonymous):

\[a=\log_{4} 16\]Take anti-log of both sides....

OpenStudy (anonymous):

\[a^4=16\]\[2^4=16\]a-2

OpenStudy (anonymous):

a=2*

OpenStudy (anonymous):

It's a log base 4, you would exponentiate both sides to the base of 4. I'm not familiar with a log base 4 anti-log. They use these numbers because it calculates either way. They want you to know the rule.

OpenStudy (anonymous):

\[\log_{b}x=y <=>b ^{y}=x \]